a ball is projected vertically upwards from a height of 2m above the ground level.its initial speed is 8m/s.find the maximum height reached by the ball above the ground


1
Expert's answer
2021-05-27T10:38:05-0400

The maximum height is the 2 m plus the height reached by the ball:


"H_\text{max}=h+h_\text{up},\\\space\\
H_\text{max}=h+\frac{v^2}{2g},\\\space\\
H_\text{max}=2+\frac{8^2}{2·9.8}=5.27\text{ m}"